OpenStudy (anonymous):

what is the area bounded by y=(cos x) e^sinx and the x axis between x=o and x=pi/2

OpenStudy (anonymous):

you need to integrate it first that will be e^sinx and than take : e^sin(pi/2)- e^sin(0) that is e-1

OpenStudy (anonymous):

e-1, or e^-1

OpenStudy (anonymous):

e-1

OpenStudy (anonymous):

as sin(pi/2)=1 sin0=0 and e^0=1

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

intergrate \[\int\limits_{?}^{?}\]3e^7x dx.

OpenStudy (anonymous):

3/7 e^7x +C

OpenStudy (anonymous):

express log13 in therms of nautural logrithms 7

OpenStudy (anonymous):

wooah sorry log_7 13

OpenStudy (anonymous):

\[\log_{7} 13\] ??

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

log7(13)=ln(13)/ln7

OpenStudy (anonymous):

I had to search that up, did not remember :-)

OpenStudy (anonymous):

okay so all i have left to do is word problems that are suber long :/

OpenStudy (anonymous):

good luck for it

OpenStudy (anonymous):

thanks